# Knowledge Graph Pruning for Recommendation

### v1

@@ 3.1.2. Holistic View
Notably, we still limit the range of the holistic importance score between 0 and 1. By leveraging the advantages of the recommender system, the holistic scores possess the capacity to represent the significance of the entities. In other words, entities with low importance scores are ineffective at capturing valuable information for the recommendation task, whereas informative entities with high importance scores are the key drivers in enhancing the performance of the recommendation system.
@@ 3.1.3. Aggregated Importance Scores
where $\gamma$ is a trade-off hyper-parameter. In this way, $\boldsymbol{\hat{s}}$ can be applied for all entities so as to discover the unnecessary or uninformative entities that are expected to be pruned from the KG. Subsequently, we design a transformation function $\mathcal{T}$ to to measure each triplet $(h,r,t)$ according to the acquired $\boldsymbol{\hat{s}}$:
@@ 3.2. Importance-Aware Graph Neural Network
In this part, we aim to establish the connection between the recommendation system and the evaluator module. In detail, we inject the final importance scores into the message propagation process of formal KGRS models and make predictions on the unobserved user-item interactions. Previous Graph-based approaches have shown that the neighborhood-based aggregation scheme is a promising method to capture the user preference with the help of KG, such as KGIN (Wang et al., 2021a) and KGRec (Yang et al., 2023). Specifically, after broadcasting the importance scores of entities to the corresponding triplets, we emphasize that the key point turns to incorporating the aggregated importance scores into the message propagation process through a dedicated graph neural network and learning the refined representations of users, items, and entities on the CKG. To be specific, we illustrate knowledge graph propagation and user-item graph propagation in detail.
@@ 3.2.2. User-Item Graph Propagation
Here $CF$ can be any collaborative filtering method to update user embeddings. $\boldsymbol{e}_{u}^{(l-1)}$ and $\boldsymbol{e}_{u}^{(l)}$ are the user embeddings of previous layer $l-1$ and current layer $l$, and $\boldsymbol{e}_{i}^{(l-1)}$ is the item embedding of previous layer $l-1$. Specifically, we simply adopt the following aggregation method to update user embeddings:
@@ 3.3. Pruned Knowledge Graph Generation
In this section, a final pruned knowledge graph $\tilde{G}$ will be generated for the downstream tasks based on $\boldsymbol{\hat{s}}(h,r,t)$. In fact, during the training process, the candidate mask $\boldsymbol{\hat{s}}^{k}$ can be generated in each epoch. At last, the objective mask vector $\boldsymbol{\tilde{s}}$ can be presented as:
@@
where $K$ is number of training epochs, $AGG$ is the function to aggregate all learned mask vectors of each epoch. Additionally, to map the scores to $\{0,1\}$, we take $\boldsymbol{s}=\boldsymbol{1}(\boldsymbol{\tilde{s}}\geq\tau)$ to mask edges in knowledge graph $G$, here $\boldsymbol{1}$ is the indicator function and $\tau$ is a threshold. Consequently, the pruned KG can be constructed as:
(15) $\tilde{G}=G\odot\boldsymbol{s}$
@@ 4.1. Experimental Settings
To ensure the data quality, we use the 10-core setting to guarantee that each user or item has at least 10 interactions. Similarly, the infrequent entities (less than 10 triplets) and relations (less than 50 triplets) are filtered out for the KG part of both datasets. It is worth noting that we utilize the same processed datasets released by KGIN (Wang et al., 2021a) and KGAT (Wang et al., 2019a), respectively.
@@
In the evaluation phase, we follow the previous studies and adopt the all-ranking strategy for convincing experiments. To measure the performance, two representative evaluation metrics for top-K recommendation, i.e., $Recall@K$ and $NDCG@K$ (Normalized Discounted Cumulative Gain) are adopted for all experiment settings. Note we set $K=20$ by default.
@@
Table 2. The experimental comparison among a wide range of knowledge graph pruning methods for three benchmarking datasets. Note that KGIN and KGRec are selected KG-based recommendation models as the backbone. R@20 and N@20 stand for Recall@20 and NDCG@20, respectively. The best results are in bold and the secondary best results are underlined.
@@
Dataset Backbone Ratio Metric Pop KGRS-IMP RotatE KGAT CAGED KGist KGTrimmer
Amazon-book KGIN 90% R@20 0.1234 0.1647 0.1573 0.1621 0.1132 0.1196 0.1667
N@20 0.0648 0.0890 0.0850 0.0871 0.0587 0.0636 0.0898
95% R@20 0.1285 0.1460 0.1560 0.1620 0.1204 0.1198 0.1662
N@20 0.0683 0.0775 0.0838 0.0871 0.0627 0.0637 0.0890
98% R@20 0.1270 0.1246 0.1423 0.1508 0.1260 0.1219 0.1615
N@20 0.0670 0.0648 0.0753 0.0803 0.0659 0.0636 0.0861
KGRec 90% R@20 0.1298 0.1549 0.1555 0.1561 0.1113 0.1141 0.1620
N@20 0.0685 0.0821 0.0837 0.0834 0.0573 0.0603 0.0861
95% R@20 0.1236 0.1443 0.1542 0.1563 0.1189 0.1138 0.1626
N@20 0.0650 0.0757 0.0821 0.0836 0.0618 0.0600 0.0860
98% R@20 0.1240 0.1225 0.1410 0.1460 0.1179 0.1120 0.1557
N@20 0.0647 0.0640 0.0745 0.0774 0.0612 0.0573 0.0826
Last-FM KGIN 30% R@20 0.0965 0.0964 0.0965 0.0922 0.0955 0.0932 0.0966
N@20 0.0843 0.0849 0.0835 0.0799 0.0825 0.0815 0.0838
70% R@20 0.0848 0.0857 0.0899 0.0830 0.0893 0.0866 0.0938
N@20 0.0738 0.0749 0.0782 0.0714 0.0773 0.0753 0.0824
90% R@20 0.0730 0.0791 0.0790 0.0738 0.0803 0.0798 0.0889
N@20 0.0626 0.0703 0.0704 0.0647 0.0689 0.0687 0.0789
KGRec 30% R@20 0.0927 0.0926 0.0922 0.0884 0.0906 0.0905 0.0941
N@20 0.0811 0.0809 0.0799 0.0764 0.0778 0.0785 0.0809
70% R@20 0.0838 0.0788 0.0863 0.0790 0.0837 0.0843 0.0909
N@20 0.0718 0.0695 0.0742 0.0681 0.0722 0.0728 0.0793
90% R@20 0.0775 0.0798 0.0753 0.0780 0.0805 0.0781 0.0870
N@20 0.0654 0.0690 0.0655 0.0674 0.0688 0.0669 0.0762
Alibaba-Fashion KGIN 30% R@20 0.1163 0.1171 0.1143 0.1070 0.1107 0.1169 0.1179
N@20 0.0724 0.0731 0.0712 0.0648 0.0686 0.0729 0.0738
70% R@20 0.1112 0.1062 0.1098 0.1018 0.0967 0.1051 0.1158
N@20 0.0690 0.0643 0.0679 0.0615 0.0588 0.0644 0.0725
90% R@20 0.1028 0.0975 0.1025 0.0970 0.0936 0.0965 0.1112
N@20 0.0627 0.0590 0.0627 0.0583 0.0562 0.0579 0.0694
KGRec 30% R@20 0.1194 0.1199 0.1176 0.1184 0.1141 0.1185 0.1199
N@20 0.0749 0.0749 0.0734 0.0741 0.0707 0.0741 0.0752
70% R@20 0.1114 0.1072 0.1101 0.1016 0.1000 0.1072 0.1165
N@20 0.0692 0.0654 0.0679 0.0614 0.0611 0.0658 0.0729
90% R@20 0.1027 0.0994 0.1010 0.0978 0.0925 0.0985 0.1104
N@20 0.0625 0.0605 0.0611 0.0590 0.0552 0.0589 0.0690
@@
  • KGIN (Wang et al., 2021a): This work considers user-item interactions at the finer granularity of intents that are a combination of KG relations for better model capacity and interpretability for the recommendation. To this end, the proposed KGIN utilizes a relational path-aware aggregation module to identify the influences of different intents.
@@
  • RotatE(Sun et al., 2019): It is a link prediction method in knowledge graph embedding that defines each relation as a rotation from the head entity to the tail entity in the complex vector space. Here, the edges with lower prediction scores will be removed in our task.
  • KGAT(Wang et al., 2019a): This work explicitly models the high-order connectivities in KG and employs an attention mechanism to tell the importance of neighbors. To adapt our task, we directly sort the learned edge scores and prune the lower ones, resulting in a pruned KG.
  • KGist(Belth et al., 2020): The state-of-the-art KG summarization method innovatively adopts the Minimum Description Length principle (MDL) to inductively compress the original KG.
  • CAGED(Zhang et al., 2022): It is a state-of-the-art error detection method in KG that predicts the scores of triplets by multi-view contrastive learning. And we finally remove the anomaly entities with connected edges to get the pruned KG.
@@
We implement our KGTrimmer in PyTorch. For a fair comparison, we fix the embedding size to 64 for all methods. We optimize the model parameters with Adam optimizer. Learning rate, regularization coefficient and dropout rate adhere to the settings of previous works(Wang et al., 2021a). In detail, the learning rate is searched amongst $\{1e^{-5},1e^{-4},1e^{-3}\}$, the regularization coefficient is tuned in $\{1e^{-6},1e^{-5},1e^{-4},1e^{-3}\}$. As for the GNN structure, the depth of GNN layers is selected from $\{1,2,3\}$, We leverage dropout operations to avoid over-fitting, where the node dropout and message dropout ratios are both tuned amongst $\{0.1,0.2,...,0.7\}$ separately. Furthermore, we choose at most $k\in\{20,50,100,200\}$ users for each entity to construct user-entity matrix. As for the transform function $\mathcal{T}$, we select from the identity function, average function and multiplicative function. Moreover, we adopt the early stopping strategy by performing premature stopping if the performance does not increase for 100 successive epochs.

### v2

@@ 4.3.2. Holistic View
To ensure comparability with the collective score, we still limit the range of the holistic importance score between 0 and 1. We have also considered other approaches, such as obtaining specific scores by passing the entity embeddings through a Multi-Layer Perceptron (MLP). Undoubtedly, MLP possesses a more powerful capability for representation, and thus is commonly used for feature extraction and fusion. However, we still decided to forgo the use of MLP for the following reasons:(1) Effectiveness. In the context of overall framework, the embeddings evolve continuously during the training phase, making it challenging for an MLP to capture their characteristic information effectively. (2) Efficiency. Compared to directly assigning scores to each entity, MLP contains more parameters and requires additional computational resources to support both training and inference.
By leveraging the advantages of the recommender system, the holistic scores possess the capacity to represent the significance of the entities. In other words, entities with low importance scores are ineffective at capturing valuable information for the recommendation task, whereas informative entities with high importance scores are the key drivers in enhancing the performance of the recommendation system.
@@ 4.3.3. Aggregated Importance Scores
where $\gamma$ is a trade-off hyper-parameter. In this way, $\boldsymbol{\hat{s}}$ can be applied for all entities so as to discover the unnecessary or uninformative entities that are expected to be pruned from the KG. Subsequently, we design a transformation function $\mathcal{T}$ to to measure each triplet $(h,r,t)$ according to the $\boldsymbol{\hat{s}}$:
@@ 4.4. Importance-Aware Graph Neural Network
In this part, we aim to establish the connection between the recommendation system and the evaluator module, i.e. importance-aware predictor $P_{\theta}(i|u,\mathcal{G},\hat{\boldsymbol{s}})$. In detail, we inject the final importance scores into the message propagation process of formal KGRS models and make predictions on the unobserved user-item interactions. Previous Graph-based approaches have shown that the neighborhood-based aggregation scheme is a promising method to capture the user preference with the help of KG, such as KGIN (Wang et al., 2021b) and KGRec (Yang et al., 2023). Specifically, after broadcasting the importance scores of entities to the corresponding triplets, we emphasize that the key point turns to incorporating the aggregated importance scores into the message propagation process through a dedicated graph neural network and learning the refined representations of users, items, and entities on the CKG. To be specific, we illustrate knowledge graph propagation and user-item graph propagation in detail.
@@ 4.4.2. User-Item Graph Propagation
Here $CF$ can be any collaborative filtering method to update user embeddings. $\boldsymbol{e}_{u}^{(l-1)}$ and $\boldsymbol{e}_{u}^{(l)}$ are the user embeddings of previous layer $l-1$ and current layer $l$, and $\boldsymbol{e}_{i}^{(l-1)}$ is the item embedding of previous layer $l-1$. Specifically, we simply adopt the following method to update user embeddings:
@@
The key distinctions between KG propagation and collaborative graph propagation parts can be summarized as follows: (1) A knowledge graph contains a variety of relationships, necessitating consideration of the effects these relations impart during propagation. In contrast, the collaborative graph is characterized by a singular type of relationship: interaction. This eliminates the necessity for explicit modeling of this interaction to avoid an increase in model complexity. (2) To filter out irrelevant information, message propagation within the KG is weighted by a importance score $\hat{\boldsymbol{s}}(v,r,t)$, which we derived from dual-view evaluator. However, for the CF side, we consider each interaction inherently valuable to the recommendation system and therefore do not perform any filtration. Thus each edge is inherently weighted with a score of 1 to demonstrate the significance of the interaction.
@@ 4.5. Pruned Knowledge Graph Generation
In this section, a final pruned knowledge graph $\tilde{G}$ will be generated for the downstream tasks based on $\boldsymbol{\hat{s}}(h,r,t)$. To achieve this goal, we still be urgent to tackle two challenges: 1) How to aggregate scores from different training stages. In the early stages of training, the model focus more on generic knowledge to fast adapt to the specific task. While in the later stages of training, the model will lay stress on refining details to better fit the training set. As a result, the importance score $\hat{\boldsymbol{s}}$ will evolve as training stages change. Thus, it is worth exploring to find an appropriate to aggregate these scores for the final pruning. 2) How to binarize the final estimated mask vector. Given that our ultimate goal is to produce a pruned knowledge graph, we need to obtain a binarized mask vector to decide which triplets should be retained and which should be discarded. Therefore, to bridge the gap between the obtained scores and the desired mask binary vector, a binarization technique is supposed to be employed.
Masked Score Aggregation. In this part, our objective is to identify an effective aggregation approach to pinpoint the least valuable information. In fact, during the training process, the candidate mask $\boldsymbol{\hat{s}}^{k}$ can be generated in each epoch. Without loss of generality, the objective mask vector $\boldsymbol{\tilde{s}}$ can be presented as:
@@
where $K$ is number of training epochs, $AGG$ is the function to aggregate all learned mask vectors of each epoch.
Masked Score Binarization. As previously discussed, it’s essential to transform the scores into a binary format to generate the ultimate pruned graph. The most straightforward strategy involves employing a threshold-based filtering technique, where setting a threshold value $\tau$, allows for the retention of triplets with scores above $\tau$ and the elimination of those below, that is:
(19) $s_{i}=\begin{cases}1,&\text{if }\tilde{s}_{i}\geq\tau\text{ ,}\\ 0,&\text{otherwise}\end{cases}$
where $i\in\{1,2,\ldots,|\mathcal{G}|\}$ represents the index of triplets. Alternatively, we also implement a percentile-based selection method to retain a certain amount of triplets. In this way, this would involve sorting all scores of triplets and only keeping the desired top $X$ percent of the triplets. Selecting an appropriate threshold facilitates the exclusion of irrelevant nodes by the system.
Indeed, both methods have their strengths and weaknesses. As for the threshold filtering approach, its advantage lies in removing nodes deemed irrelevant, simultaneously simplifying the dataset while retaining higher-scoring triplets. This method is well-suited for practical application scenarios. For the percentile-based method, it allows for the specification of a pruning ratio, making it more convenient to compare with other baselines. Therefore, to demonstrate the superiority of our method, we have adopted the percentile-based method. More details about the relation between these two methods will be included in the section 5.2.5.
Consequently, the pruned KG can be constructed as:
(20) $\tilde{\mathcal{G}}=\mathcal{G}\odot\boldsymbol{s}$
where $\boldsymbol{s}$ is the estimated binary mask vector. Finally, our algorithm is summarized in Algorithm 1.
Data: users $\mathcal{U}$, items $\mathcal{I}$, user-item interactions $Y$, and knowledge graph $\mathcal{G}$;
Input: batch size $B$, number of layers $L$, balance coefficient $\gamma$, $k$ users per entity;
Output: pruned knowledge graph $\tilde{\mathcal{G}}$;
Construct the user-entity matrix ${Q}$ given the $k$;
Initialize learnable embeddings $\boldsymbol{e}_{u}^{(0)}$, $\boldsymbol{e}_{v}^{(0)}$, $\boldsymbol{e}_{r}$ via Xavier Initialization;
Assign the learnable parameters $\boldsymbol{e}^{h}$ to each node and initialize as 1;
Set empty list to $\mathcal{S}$ ;
while not convergent do
       Sample $B$ user-item pairs from $Y$;
       Construct negative samples $Y^{-}$ for each interaction and build training data to feed the KGPR model;
       Lookup the corresponding entities $v$ from $Q$ according to $u$ that appears in the batch;
       $\boldsymbol{s}^{c}=sim(\boldsymbol{e}_{u}^{(0)},\boldsymbol{e}_{v}^{\prime})$ ;
       Gather the holistic scores $\boldsymbol{s}^{h}$;
       $\hat{\boldsymbol{s}}=\gamma\boldsymbol{s}^{c}+(1-\gamma)\boldsymbol{s}^{h}$;
       Append $\hat{\boldsymbol{s}}$ to $\mathcal{S}$;
       for $l\in\left(1,2,\ldots,L\right)$ do
             Calculate the $l$-order representations of $e_{v}$;
             $\boldsymbol{e}_{v}^{(l)}=KGE(\boldsymbol{e}_{v}^{(l-1)},\boldsymbol{e}_{r},\hat{\boldsymbol{s}})$;
             Calculate the $l$-order representation of $e_{u}$;
             $\boldsymbol{e}_{u}^{(l)}=CF(\boldsymbol{e}_{u}^{(l-1)},\boldsymbol{e}_{i}^{(l-1)})$;
       end for
      Calculate the final Embeddings $\boldsymbol{e}_{u}$ and $\boldsymbol{e}_{i}$;
      Calculate the BPR loss $\mathcal{L}$;
      Update $\Theta=\left\{\boldsymbol{e}_{u}^{(0)},\boldsymbol{e}_{v}^{(0)},\boldsymbol{s}^{h}\right\}$ according to gradient descent;
end while
$\tilde{\boldsymbol{s}}=AGG(\mathcal{S})$ ;
Generate the final pruned Graph $\tilde{\mathcal{G}}$ by percentile-based selection;
return $\tilde{\mathcal{G}}$ ;
Algorithm 1 Algorithm of KGPR
@@ 5.1. Experimental Settings
To ensure the data quality, we use the 10-core setting to guarantee that each user or item has at least 10 interactions. Similarly, the infrequent entities (less than 10 triplets) and relations (less than 50 triplets) are filtered out for the KG part of both datasets. It is worth noting that we utilize the same processed datasets released by KGIN (Wang et al., 2021b) and KGAT (Wang et al., 2019a), respectively.
@@
In the evaluation phase, we follow the previous studies and adopt the all-ranking strategy for convincing experiments. To measure the performance, two representative evaluation metrics for top-K recommendation, i.e., $Recall@K$ and $NDCG@K$ (Normalized Discounted Cumulative Gain) are adopted for all experiment settings. Note we set $K\in\{10,20\}$, as small $K$ may encounter the instability of performance, while large $K$ may mismatch the nature of recommendation system that user prefer to clicking the top-K items.
@@
Table 2. The experimental comparison among a wide range of knowledge graph pruning methods for three benchmarking datasets. Note that KGIN is selected KG-based recommendation model as the backbone. R@K and N@K stand for Recall@K and NDCG@K, respectively. The best results are in bold and the secondary best results are underlined.
@@
Dataset Ratio Metric Pop KGRS CAGED KGist MoG KRDN KGTrimmer
Amazon-book 90% R@10 0.0844 0.1133 0.0736 0.0805 0.0796 0.0819 0.1155
R@20 0.1272 0.1641 0.1132 0.1196 0.1217 0.1217 0.1667
N@10 0.0544 0.0729 0.0467 0.0515 0.0506 0.0528 0.0741
N@20 0.0675 0.0885 0.0587 0.0636 0.0635 0.0650 0.0898
95% R@10 0.0860 0.0987 0.0803 0.0805 0.0786 0.0787 0.1152
R@20 0.1285 0.1464 0.1204 0.1198 0.1216 0.1177 0.1662
N@10 0.0553 0.0625 0.0504 0.0512 0.0498 0.0503 0.0733
N@20 0.0683 0.0772 0.0627 0.0637 0.0630 0.0622 0.0890
98% R@10 0.0847 0.0851 0.0833 0.0801 0.0824 0.0805 0.1099
R@20 0.1270 0.1276 0.1260 0.1219 0.1266 0.1210 0.1615
N@10 0.0541 0.0540 0.0527 0.0508 0.0526 0.0511 0.0703
N@20 0.0670 0.0671 0.0659 0.0636 0.0660 0.0635 0.0861
Last-FM 30% R@10 0.0688 0.0685 0.0665 0.0656 0.0647 0.0672 0.0690
R@20 0.0959 0.0963 0.0947 0.0925 0.0922 0.0949 0.0966
N@10 0.0762 0.0762 0.0738 0.0728 0.0725 0.0741 0.0758
N@20 0.0838 0.0841 0.0818 0.0803 0.0804 0.0819 0.0838
70% R@10 0.0603 0.0589 0.0615 0.0597 0.0560 0.0612 0.0667
R@20 0.0856 0.0844 0.0885 0.0854 0.0814 0.0873 0.0936
N@10 0.0675 0.0676 0.0692 0.0673 0.0635 0.0685 0.0746
N@20 0.0743 0.0744 0.0768 0.0746 0.0705 0.0758 0.0819
90% R@10 0.0545 0.0573 0.0562 0.0561 0.0527 0.0553 0.0627
R@20 0.0785 0.0818 0.0826 0.0804 0.0766 0.0803 0.0889
N@10 0.0607 0.0649 0.0638 0.0630 0.0603 0.0627 0.0720
N@20 0.0675 0.0714 0.0711 0.0698 0.0668 0.0696 0.0789
Alibaba-Fashion 30% R@10 0.0761 0.0769 0.0726 0.0761 0.0708 0.0736 0.0773
R@20 0.1161 0.1169 0.1111 0.1163 0.1090 0.1123 0.1179
N@10 0.0579 0.0589 0.0551 0.0583 0.0538 0.0562 0.0592
N@20 0.0723 0.0732 0.0690 0.0728 0.0675 0.0691 0.0738
70% R@10 0.0725 0.0690 0.0604 0.0667 0.0596 0.0652 0.0761
R@20 0.1112 0.1070 0.0951 0.1039 0.0943 0.1017 0.1154
N@10 0.0548 0.0515 0.0452 0.0501 0.0442 0.0488 0.0582
N@20 0.0687 0.0652 0.0577 0.0634 0.0567 0.0619 0.0723
90% R@10 0.0656 0.0614 0.0592 0.0604 0.0604 0.0623 0.0725
R@20 0.1018 0.0966 0.0930 0.0959 0.0945 0.0970 0.1114
N@10 0.0493 0.0457 0.0439 0.0446 0.0445 0.0465 0.0553
N@20 0.0623 0.0583 0.0560 0.0573 0.0568 0.0589 0.0692
Table 3. The experimental comparison among a wide range of knowledge graph pruning methods for three benchmarking datasets. Note that KGRec is selected KG-based recommendation model as the backbone. R@K and N@K stand for Recall@K and NDCG@K, respectively. The best results are in bold and the secondary best results are underlined.
Dataset Ratio Metric Pop KGRS CAGED KGist MoG KRDN KGTrimmer
Amazon-book 90% R@10 0.0861 0.1086 0.0750 0.0767 0.0799 0.0748 0.1111
R@20 0.1284 0.1569 0.1138 0.1153 0.1235 0.1136 0.1620
N@10 0.0553 0.0695 0.0473 0.0489 0.0504 0.0471 0.0710
N@20 0.0682 0.0845 0.0592 0.0607 0.0636 0.0589 0.0867
95% R@10 0.0836 0.0947 0.0793 0.0776 0.0772 0.0700 0.1110
R@20 0.1251 0.1421 0.1193 0.1172 0.1195 0.1067 0.1625
N@10 0.0535 0.0598 0.0500 0.0498 0.0485 0.0441 0.0701
N@20 0.0662 0.0743 0.0623 0.0619 0.0615 0.0554 0.0860
98% R@10 0.0865 0.0847 0.0840 0.0785 0.0774 0.0730 0.1053
R@20 0.1293 0.1265 0.1265 0.1185 0.1199 0.1118 0.1557
N@10 0.0550 0.0535 0.0528 0.0498 0.0482 0.0459 0.0671
N@20 0.0680 0.0663 0.0658 0.0621 0.0613 0.0578 0.0826
Last-FM 30% R@10 0.0655 0.0659 0.0639 0.0640 0.0625 0.0636 0.0653
R@20 0.0938 0.0942 0.0919 0.0912 0.0889 0.0914 0.0927
N@10 0.0732 0.0728 0.0712 0.0708 0.0703 0.0710 0.0727
N@20 0.0812 0.0810 0.0789 0.0786 0.0776 0.0789 0.0803
70% R@10 0.0581 0.0565 0.0579 0.0560 0.0552 0.0600 0.0638
R@20 0.0836 0.0820 0.0844 0.0812 0.0806 0.0851 0.0915
N@10 0.0643 0.0646 0.0654 0.0638 0.0621 0.0670 0.0717
N@20 0.0715 0.0714 0.0727 0.0708 0.0693 0.0740 0.0793
90% R@10 0.0535 0.0548 0.0550 0.0539 0.0507 0.0540 0.0605
R@20 0.0784 0.0786 0.0815 0.0784 0.0745 0.0793 0.0875
N@10 0.0582 0.0621 0.0615 0.0606 0.0571 0.0608 0.0692
N@20 0.0658 0.0683 0.0691 0.0675 0.0638 0.0679 0.0764
Alibaba-Fashion 30% R@10 0.0787 0.0787 0.0747 0.0781 0.0736 0.0758 0.0794
R@20 0.1195 0.1196 0.1135 0.1180 0.1125 0.1157 0.1198
N@10 0.0603 0.0603 0.0566 0.0595 0.0560 0.0579 0.0609
N@20 0.0749 0.0749 0.0706 0.0739 0.0700 0.0722 0.0754
70% R@10 0.0731 0.0688 0.0636 0.0695 0.0624 0.0668 0.0767
R@20 0.1117 0.1070 0.0993 0.1069 0.0978 0.1039 0.1165
N@10 0.0558 0.0517 0.0478 0.0525 0.0468 0.0503 0.0585
N@20 0.0697 0.0654 0.0606 0.0659 0.0595 0.0637 0.0729
90% R@10 0.0667 0.0634 0.0576 0.0625 0.0607 0.0624 0.0728
R@20 0.1033 0.0991 0.1006 0.0981 0.0956 0.0983 0.1114
N@10 0.0500 0.0474 0.0424 0.0460 0.0448 0.0463 0.0556
N@20 0.0632 0.0603 0.0546 0.0588 0.0573 0.0592 0.0690
@@
  • KGIN (Wang et al., 2021b): This work considers user-item interactions at the finer granularity of intents that are a combination of KG relations for better model capacity and interpretability for the recommendation. To this end, the proposed KGIN utilizes a relational path-aware aggregation module to identify the influences of different intents.
@@
  • KGist(Belth et al., 2020): The promising KG summarization method innovatively adopts the Minimum Description Length principle (MDL) to inductively compress the original KG.
  • CAGED(Zhang et al., 2022): It is a popular error detection method in KG that predicts the scores of triplets by multi-view contrastive learning. And we finally remove the anomaly entities with connected edges to get the pruned KG.
  • MoG(Zhang et al., 2024a): This work exerts large effort on graph pruning task, which utilizes multiple sparsifier experts with unique sparsity level and pruning criteria for each node.
  • KRDN(Zhu et al., 2023): It is the state-of-the-art denoising knowledge-aware recommendation scheme that tries to distill the high-quality KG triplets and prune the noisy edges.
@@
We implement our KGTrimmer in PyTorch. For a fair comparison, we utilize the Xavier initializer to initialize the embeddings, where the embedding size is fixed to 64 for all methods. We optimize the model parameters with Adam optimizer, where the batch size is set to 4096. The grid search strategy is applied to search hyper-parameters for reaching the optimal performance. Learning rate, regularization coefficient and dropout rate adhere to the settings of previous works(Wang et al., 2021b). In detail, the learning rate is searched amongst $\{1e^{-5},1e^{-4},1e^{-3}\}$, the regularization coefficient is tuned in $\{1e^{-6},1e^{-5},1e^{-4},1e^{-3}\}$. As for the GNN structure, the depth of GNN layers is selected from $\{1,2,3\}$, We leverage dropout operations to avoid over-fitting, where the node dropout and message dropout ratios are both tuned amongst $\{0.1,0.2,...,0.7\}$ separately. Furthermore, we choose at most $k\in\{20,50,100,200\}$ users for each entity to construct user-entity matrix. As for the transform function $\mathcal{T}$, we select from the identity function, average function and multiplicative function. Moreover, we adopt the early stopping strategy by performing premature stopping if the performance does not increase for 100 successive epochs.